# The impact of $\gamma N$ and $\gamma^{*}N$ interactions on our understanding of nucleon excitations

### v1

@@ 1   Introduction
Quantum Chromodynamics (QCD), the theory describing the interaction between colored quarks and colored gluons [1], encompasses a vast range of physics. Over time, it explains how quarks and gluons formed a quark-gluon plasma (QGP) soon after the end of the Universe’s inflationary phase and the subsequent formation of mesons and baryons approximately 10 $\mu$s later. QCD is essential for understanding stellar evolution and the structure of protons and neutrons. Moreover, it plays a crucial role in various areas of modern physics, such as calculations of the “background" for Higgs particle production at the LHC at multi-TeV energies and determining the hadronic contributions to the muon’s anomalous magnetic moment [2].
Across this broad spectrum, the QCD coupling constant $\alpha_{s}$ varies significantly. At high momentum transfers, $\alpha_{s}$ is small, allowing perturbative methods to be applied for computing physical quantities. However, at low momentum transfers, $\alpha_{s}$ increases, necessitating non-perturbative calculations. Despite these challenges, a comprehensive view of the coupling constant is illuminating (see Fig. 1.1). Perturbative determinations are compiled by the Particle Data Group [3], while non-perturbative results with references to the relevant calculations can be found in Ref. [4].
Quarks confined within hadrons experience very low momentum transfers, making their spectrum difficult to calculate and understand. Gell-Mann [5] and Zweig [6] proposed to interprete mesons as bound states of a quark and an antiquark, and baryons as bound states of three quarks. The meson spectrum already presents a challenge — so why should we focus on baryons?
“Why $N^{*}$’s?” This was the question with which Nathan Isgur began his talk at $N^{*}$2000 [7], held at the Thomas Jefferson National Accelerator Facility in Newport News, VA, one year before his untimely passing. He provided three answers: First, “Nucleons are the building blocks of our world. As such, they must be central to any discussion of why the world we experience has the characteristics it does." Second, nucleons “are the simplest system in which the quintessential non-Abelian character of QCD is manifest. After all, there are $N_{c}$ quarks in a proton because there are $N_{c}$ colors.” Third, “baryons are sufficiently complex to reveal physics that remains hidden to us in the mesons." Gell-Mann and Zweig did not fully develop their quark model based on mesons, whose simpler structure allowed for alternative interpretations.
Figure 1.1: (color online) The strong interaction coupling constant, covering the non-perturbative to perturbative regions. References to the data can be found in [4].
@@
Several excellent reviews on nucleon excitations have been written. In 1982, Hey and Kelly [8] reviewed symmetries and quark model calculations, comparing experimental results from $\pi N$ interactions with theoretical predictions. First results from the new electron accelerator facilities were discussed by Burkert and Lee [9], where the formalism for electromagnetic meson production was laid out and theoretical models discussed. Early photoproduction experiments were summarized by Klempt and Richard [10], while Crede and Roberts [11] discussed advancements in data analysis and new observables. Ireland, Pasyuk, and Strakovsky [12] reviewed photon beam experiments and compiled an extensive set of experimental results with comprehensive references. Eichmann delved into baryon theory, including quark-model wave functions and advancements in solving the three-body Bethe-Salpeter equation [13]. Thiel, Afzal, and Wunderlich [14] provided an extensive review of baryon spectra, spin formalisms, and the state of “complete experiments," summarizing major data sets and partial-wave analyses. Aznauryan and Burkert [15, 16] and Carman, Jo, and Mokeev [17] reviewed progress in understanding nucleon resonance structures through meson electroproduction data, exploring the interplay between a core of three dressed quarks and the surrounding meson-baryon cloud. Ramalho and Peña [18] wrote a comprehensive review on recent experimental and theoretical achievements in our understanding of the electromagnetic structure of nucleons and nucleon excitations. The authors of Ref. [19] focus on nucleon resonance electroexcitation and the emergence of hadron mass.
@@
In Section 4, approximations consistent with QCD are reviewed. At low energies, QCD can be approximated by effective field theories in which hadrons are the elementary building blocks. We compare the resonance spectrum to predictions of different quark models, starting from so-called constituent quarks with large constituent-quark masses. The strong interaction constant saturates at low $Q^{2}$ (see Fig.1.1), a fact exploited in AdS/QCD, which successfully predicts the full excitation spectrum of the nucleon, provided the AdS/QCD mass formula is given a new interpretation. The baryon spectrum can also be calculated using functional methods, which provide insights into QCD’s non-perturbative phenomena such as the generation of dynamical mass. Significant progress has been achieved in recent years in calculating hadron masses in lattice QCD. While the baryon ground-state masses have been calculated using quark masses close to or at their “true" values, studying the excitation spectrum is still an ongoing endeavor.
@@ 2.1.1   Pion-induced reactions
For many years, most of our knowledge about the excited states of the nucleon was based on analyses of $\pi N$ elastic scattering conducted by three major groups: the Karlsruhe-Helsinki (KH) group, the Carnegie-Mellon-Berkeley (CMB) group, and the SAID group at George Washington University. Early data on $\pi N$ elastic scattering were compiled in [20], with results reported in [21, 22, 23, 24, 25, 26, 27].
@@
The CMB group emphasized the importance of performing a rigorous statistical analysis. They constructed a new dataset with predefined energy bins and solid angles [28]. This method, termed “amalgamation", accounted for normalization errors, momentum calibration errors, momentum resolution, electromagnetic corrections, threshold structures, and inconsistencies among datasets. The full covariance matrix of the amalgamated data was calculated, incorporating both statistical and systematic errors.
@@
The SAID group has continuously updated the database. References to the data are available at Ref. [29].
@@
Only a few elastic $\pi N$ scattering data have been reported since 1990. Notably, spin rotation parameter measurements conducted at ITEP, Moscow, made significant contributions [30, 31, 32, 30, 33].
@@
Inelastic pion-induced reactions have also contributed to our understanding. Pioneering work by Longacre and collaborators [34, 35, 36, 37], as well as Novoseller [38, 39], included data on inelastic channels and reported on $N^{*}$ and $\Delta^{*}$ couplings to $N\pi\pi$ final states. Most data on inelastic pion scattering originate from bubble-chamber experiments, with analyses conducted by the Kent group. References to these data and their results are provided in [40, 41]. More recent counter experiments have focused on two-pion production in $\pi N$ scattering. At BNL, the reaction $\pi^{-}p\to\pi^{0}\pi^{0}p$ was measured in the beam momentum range 270–750 MeV using the Crystal Ball multiphoton spectrometer [42]. The HADES Collaboration investigated $\pi^{-}p\to\pi^{+}\pi^{-}p$ at four pion beam momenta: 650, 685, 733, and 786 MeV [43].
@@
Scattering of $\pi^{+}$ on protons is particularly suited for studying $\Delta^{*}$ resonances [44], as no $N^{*}$ resonances with charge $+2$ exist. Studies of the reaction $\pi^{+}p\to K^{+}\Sigma^{+}$ [45] benefit from the ability to measure spin rotation parameters using a polarized target and by detecting $\Sigma^{+}$ polarization [46].
2.1.2   Hyperon production
@@
A comprehensive list of references to data on elastic and inelastic $\bar{K}N$ scattering in the resonance region can be found in Ref. [47].
2.1.3   Photoproduction of mesons
@@
Over the past two decades, our understanding of nucleon excitations, including $N^{*}$ and $\Delta^{*}$ states, has advanced significantly due to efforts by a large scientific community studying photon-induced reactions on H2 or D2 targets.
@@
In Bonn, the first 500 MeV electron synchrotron was constructed under the leadership of Wolfgang Paul between 1953 and 1958, with initial experiments on pion production from protons reported in [48]. A second synchrotron capable of accelerating electrons up to 2.5 GeV became operational in 1967 [49]. Shortly thereafter, a butanol target with polarized protons, developed at CERN [50], was installed in Bonn [51]. Development of polarized electron sources in Bonn has a long tradition, with the first such source using photoionization of polarized lithium atoms [52]. Later, spin-polarized electrons were produced using photoemission from Be-InGaAs/AlGaAs strained layer superlattice photocathodes, driven by circularly polarized laser light [53]. The first acceleration of polarized electrons in the 2.5 GeV electron synchrotron occurred in the mid-1980’s  [54] to study the effect of the depolarizing resonances.
@@
The synchrotron initially delivered high-intensity electron bunches with an approximately 5% duty cycle. To increase this value and make coincidence measurements feasible, the Electron Stretcher Ring ELSA, capable of operating at a maximum energy of 3.5 GeV, was constructed and became operational in 1987 [55]. Several experiments relevant for baryon spectroscopy have been conducted there.
@@
The PHOENICS experiment used a polarized frozen spin target to study pion and $\eta$ photoproduction off protons and deuterons. The detector was also used by the GDH collaboration, exploiting circularly polarized photons and longitudinally polarized protons to test the Gerasimov-Drell-Hearn sum rule [56, 57, 58, 59, 60, 61, 62]. Data on the low-energy region were taken at Mainz (see below).
@@
The ELAN experiment collected data on the four-momentum transfer dependence of the $N\to\Delta$ transition by measuring angular distributions of the pion production cross section. Preliminary results are reported in [63].
@@
The first data at ELSA were taken with the SAPHIR detector [64]. This experiment was the first counter experiment on photoproduction in the resonance region, covering the full angular range. Results were reported in Refs. [65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79].
@@
The Crystal Barrel detector [80] had previously been used at the CERN Low Energy Antiproton Ring. The Two-Arm-Photon Spectrometer [81] was dismantled at Mainz and installed in Bonn to provide better granularity for the forward hole of the Crystal Barrel detector. Data were collected on meson photoproduction off protons or neutrons, both without polarization and with various combinations of photon and target polarizations [82, 83, 84, 85, 86, 87, 88, 89, 90, 91, 92, 93, 94, 95, 96].
The first accelerator built in Mainz (1964–1966) was a 300 MeV linear electron accelerator [97] for nuclear structure experiments. In the summer of 1976, a decision was made to construct a multilevel racetrack Microtron (MAMI). A fit stage, which accelerated electrons from a 2.1 MeV (later 3.46 MeV) linear accelerator to 14 MeV, was successfully tested in 1979. The second stage, with an energy of 175 MeV, went into operation [98]. The third stage, completed in 1990, accelerated electrons up to 855 MeV and delivered an intense beam [99]. The final stage, MAMI-C, was completed in 1998 and reached an energy of 1508 MeV [100], covering the mass range up to the fourth resonance region. It provides a beam current of 100 $\mu$A within a narrow energy window. Polarized electrons, typically 80% polarized, achieve a beam current of up to 20 $\mu$A. Tagged photons hit a target, and the final state of the interaction is detected in a system composed of the Crystal Ball (CB) [101] and TAPS [81] photon spectrometers. Results were reported in Refs. [102, 103, 104, 105, 106, 107, 108, 109, 110, 111, 112, 113, 114, 115, 116, 117, 118, 119, 120, 121, 122, 123, 124, 125, 126, 127, 128, 129, 130, 131, 132, 133, 134, 135, 136, 137, 138, 139, 140, 141, 142, 143].
@@
The Thomas Jefferson National Accelerator Facility Laboratory (JLab) in Newport News, Virginia, delivered its first beam in 1994. A central topic in its wide program is pursued with the CLAS detector [144], measuring a variety of meson final states using unpolarized, linearly polarized, and circularly polarized photon beams with energies up to 5.6 GeV. Results are documented in [145, 146, 147, 148, 149, 150, 151, 152, 153, 154, 155, 156, 157, 158, 150, 159, 160, 161, 162, 163, 164, 165, 166, 167, 168, 169, 170, 171, 172, 173, 174, 175, 176, 177, 178, 179, 180, 181, 182, 183, 182, 184, 185, 186, 187, 188, 189, 190, 191, 192, 193, 194, 195, 196, 197, 198, 199, 200, 201, 202, 203, 204, 205, 206, 207, 208, 209, 210, 211, 212, 213, 214, 215, 216, 217, 218, 219, 220, 221, 222, 223].
@@
The GRAAL Collaboration scattered a beam of ultraviolet laser light off the electron beam in the straight section of the European Synchrotron Radiation Facility, thus producing narrowly collimated $\gamma$-rays with a maximum energy of 1.47 GeV. The energy distribution is flat, and the $\gamma$-rays carry the polarization of the primary photons. GRAAL results can be found in Refs. [224, 225, 226, 227, 228, 229, 230, 231, 232, 233, 234]. The BGO calorimeter is a central part of the GRAAL detector. After the end of data taking in Grenoble in 2008, the calorimeter was transferred to Bonn and became the central component of the BGO-OD experiment [235, 236] at ELSA. Results of the BGO-OD experiment at Bonn can be found in Refs. [237, 238, 239, 240, 241, 242, 243].
@@
At the 8 GeV synchrotron radiation facility Spring-8, situated in Harima Science Park, Hyōgo, an ultraviolet photon beam is backscattered off the circulating electron beam, producing collimated $\gamma$-rays up to 2.4 GeV [244, 245]. Charged particles produced in $\gamma N$ reactions in the forward direction are momentum-analyzed in a dipole magnet, with particle identification provided by time-of-flight and Cherenkov counters. The facility started operating in 1999. Physics results have been reported in Refs. [246, 247, 248, 249, 250, 251, 252, 253, 254, 255, 256, 257, 258, 259, 260, 261, 262, 263, 264, 265, 266, 267, 268, 269, 270, 271, 272, 273, 274, 275, 276, 277, 278, 279, 280, 281, 282].
2.1.4   Electroproduction

### v4

@@ 1  Introduction
[…]
Across this broad spectrum, the QCD coupling constant $\alpha_{s}$ varies significantly. At high momentum transfers, $\alpha_{s}$ is small, allowing perturbative methods to be applied for computing physical quantities. However, at low momentum transfers, $\alpha_{s}$ increases, necessitating non-perturbative calculations. Despite these challenges, a comprehensive view of the coupling constant is illuminating (see Fig. 1.1). Perturbative determinations are compiled by the Particle Data Group ParticleDataGroup:2024cfk , while non-perturbative results with references to the relevant calculations can be found in Ref. Deur:2022msf .
Quarks confined within hadrons experience very low momentum transfers, making the hadron spectrum difficult to calculate and understand. Gell-Mann Gell-Mann:1964ewy and Zweig Zweig:1964jf proposed to interpret mesons as bound states of a quark and an antiquark, and baryons as bound states of three quarks. The meson spectrum already presents a challenge — so why should we focus on baryons?
“Why $N^{*}$’s?” This was the question with which Nathan Isgur began his talk at $N^{*}$2000 Isgur:2000ad , held at the Thomas Jefferson National Accelerator Facility in Newport News, VA, one year before his untimely passing. He provided three answers: First, “Nucleons are the building blocks of our world. As such, they must be central to any discussion of why the world we experience has the characteristics it does." Second, nucleons “are the simplest system in which the quintessential non-Abelian character of QCD is manifest. After all, there are $N_{c}$ quarks in a proton because there are $N_{c}$ colors.” Third, “baryons are sufficiently complex to reveal physics that remains hidden to us in the mesons." Gell-Mann and Zweig did not fully develop their quark model based on mesons, whose simpler structure allowed for alternative interpretations.
Figure 1.1: (color online) The strong interaction coupling constant, covering the non-perturbative to perturbative regions. References to the data can be found in Deur:2022msf .
@@
Several excellent reviews on nucleon excitations have been written. In 1982, Hey and Kelly Hey:1982aj reviewed symmetries and quark model calculations, comparing experimental results from $\pi N$ interactions with theoretical predictions. First results from the new electron accelerator facilities were discussed by Burkert and Lee Burkert:2004sk , where the formalism for electromagnetic meson production was laid out and theoretical models discussed. Early photoproduction experiments were summarized by Klempt and Richard Klempt:2009pi , while Crede and Roberts Crede:2013kia discussed advancements in data analysis and new observables. Ireland, Pasyuk, and Strakovsky Ireland:2019uwn reviewed photon beam experiments and compiled an extensive set of experimental results with comprehensive references. Eichmann delved into baryon theory, including quark-model wave functions and advancements in solving the three-body Bethe-Salpeter equation Eichmann:2022zxn . Thiel, Afzal, and Wunderlich Thiel:2022xtb provided an extensive review of baryon spectra, spin formalisms, and the state of “complete experiments," summarizing major data sets and partial-wave analyses. Mai, Meißner, and Urbach reviewed the current state of the art of our understanding of the spectrum of excited strongly interacting particles Mai:2022eur . Dynamical coupled channel approaches aiming at a reliable extraction of resonance spectra and their properties from data were discussed by Döring, Haidenbauer, Mai, and Sato Doring:2025sgb . Aznauryan and Burkert Aznauryan:2011qj , Burkert:2022ioj and Carman, Jo, and Mokeev Carman:2020qmb reviewed progress in understanding nucleon resonance structures through meson electroproduction data, exploring the interplay between a core of three dressed quarks and the surrounding meson-baryon cloud. Ramalho and Peña Ramalho:2023hqd wrote a comprehensive review on recent experimental and theoretical achievements in our understanding of the electromagnetic structure of nucleons and nucleon excitations. The authors of Ref. Achenbach:2025kfx focus on nucleon resonance electroexcitation and the emergence of hadron mass.
@@
In Section 4, approximations consistent with QCD are reviewed. At low energies, QCD can be described by effective field theories in which hadrons are the elementary building blocks. We compare the resonance spectrum to predictions of different quark models, starting from so-called constituent quarks with large constituent-quark masses. The strong interaction constant saturates at low $Q^{2}$ (see Fig.1.1), a fact exploited in AdS/QCD, which successfully predicts the full excitation spectrum of the nucleon, provided the AdS/QCD mass formula is given a new interpretation. The baryon spectrum can also be calculated using functional methods, which provide insights into QCD’s non-perturbative phenomena such as the generation of dynamical mass. Significant progress has been achieved in recent years in calculating hadron masses in lattice QCD. While the baryon ground-state masses have been calculated using quark masses close to or at their “true" values, studying the excitation spectrum is still an ongoing endeavor.
@@ 2.1  Data
Pion-induced reactions: For many years, most of our knowledge about the excited states of the nucleon was based on analyses of $\pi N$ elastic scattering conducted by three major groups: the Karlsruhe-Helsinki (KH) group, the Carnegie-Mellon-Berkeley (CMB) group, and the SAID group at George Washington University. Early data on $\pi N$ elastic scattering were compiled in Hohler:1977em , with results reported in Hohler:1979yr , Koch:1980ay , Koch:1983ht , Hohler:1984ux , Koch:1985bp , Koch:1985bn , Hohler:1993lbk .
The KH group constructed $\pi N$ partial wave amplitudes constrained by a simultaneous fixed momentum transfer analysis, thus satisfying sus-u crossing symmetry. Different types of dispersion relations were imposed.
@@
The CMB group emphasized the importance of performing a rigorous statistical analysis. They constructed a new dataset with predefined energy bins and solid angles Kelly:1979uf . This method, termed “amalgamation", accounted for normalization errors, momentum calibration errors, momentum resolution, electromagnetic corrections, threshold structures, and inconsistencies among datasets. The full covariance matrix of the amalgamated data was calculated, incorporating both statistical and systematic errors.
@@
The SAID group has continuously updated the database. References to the data are available at Ref. gwdac .
@@
Only a few elastic $\pi N$ scattering data have been reported since 1990. Notably, spin rotation parameter measurements conducted at ITEP, Moscow, made significant contributions Alekseev:2005zr , Alekseev:1996gs , ITEP-PNPI:2000gqp , Alekseev:2005zr , ITEP-PNPI:2008cmv .
@@
Inelastic pion-induced reactions have also contributed to our understanding. Pioneering work by Longacre and collaborators Herndon:1974xd , Longacre:1975qj , Longacre:1976ja , Longacre:1977ga , as well as Novoseller Novoseller:1977xt , Novoseller:1977xs , included data on inelastic channels and reported on $N^{*}$ and $\Delta^{*}$ couplings to $N\pi\pi$ final states. Most data on inelastic pion scattering originate from bubble-chamber experiments, with analyses conducted by the Kent group. References to these data and their results are provided in Manley:1984zs , Manley:1992yb . More recent counter experiments have focused on two-pion production in $\pi N$ scattering. At BNL, the reaction $\pi^{-}p\to\pi^{0}\pi^{0}p$ was measured in the beam momentum range 270–750 MeV using the Crystal Ball multiphoton spectrometer CrystalBall:2004qln . The HADES Collaboration investigated $\pi^{-}p\to\pi^{+}\pi^{-}p$ at four pion beam momenta: 650, 685, 733, and 786 MeV HADES:2020kce .
@@
Scattering of $\pi^{+}$ on protons is particularly suited for studying $\Delta^{*}$ resonances Candlin:1984av , as no $N^{*}$ resonances with charge +2+2 exist. Studies of the reaction $\pi^{+}p\to K^{+}\Sigma^{+}$ Edinburgh-Rutherford-Westfield:1983qsm benefit from the ability to measure spin rotation parameters using a polarized target and by detecting $\Sigma^{+}$ polarization Candlin:1988pn .
@@
Hyperon production: A comprehensive list of references to data on elastic and inelastic $\bar{K}N$ scattering in the resonance region can be found in Ref. Matveev:2019igl .
@@
Photoproduction of mesons: Over the past two decades, our understanding of nucleon excitations, including $N^{*}$ and $\Delta^{*}$ states, has advanced significantly due to efforts by a large scientific community studying photon-induced reactions on H2 or D2 targets.
@@
In Bonn, the first 500 MeV electron synchrotron was constructed under the leadership of Wolfgang Paul between 1953 and 1958, with initial experiments on pion production from protons reported in Althoff:1963hfa . A second synchrotron capable of accelerating electrons up to 2.5 GeV became operational in 1967 Althoff:1968twa . Shortly thereafter, a butanol target with polarized protons, developed at CERN Mango:1969ww , was installed in Bonn Althoff:1971zza . Development of polarized electron sources in Bonn has a long tradition, with the first such source using photoionization of polarized lithium atoms Baum:1969es . Later, spin-polarized electrons were produced using photoemission from Be-InGaAs/AlGaAs strained layer superlattice photocathodes, driven by circularly polarized laser light Hillert:2006yb . The first acceleration of polarized electrons in the 2.5 GeV electron synchrotron occurred in the mid-1980’s  Brefeld:1984iv to study the effect of the depolarizing resonances.
@@
The synchrotron initially delivered high-intensity electron bunches with an approximately 5% duty cycle. To increase this value and make coincidence measurements feasible, the Electron Stretcher Ring ELSA, capable of operating at a maximum energy of 3.5 GeV, was constructed and became operational in 1987 Husmann:1988vg . Several experiments relevant for baryon spectroscopy have been conducted there.
@@
The PHOENICS experiment used a polarized frozen spin target to study pion and $\eta$ photoproduction off protons and deuterons. The detector was also used by the GDH collaboration, exploiting circularly polarized photons and longitudinally polarized protons to test the Gerasimov-Drell-Hearn sum rule Dutz:1996uc , Dutz:2004zz , Hoffmann-Rothe:1997jmq , Bock:1998rk , GDH:2003xhc , Naumann:2003vf , GDH:2005noz . Data on the low-energy region were taken at Mainz (see below).
@@
The ELAN experiment collected data on the four-momentum transfer dependence of the $N\to\Delta$ transition by measuring angular distributions of the pion production cross section. Preliminary results are reported in Gothe:2001yjy .
@@
The first data at ELSA were taken with the SAPHIR detector Schwille:1994vg . This experiment was the first counter experiment on photoproduction in the resonance region, covering the full angular range. Results were reported in Refs. Bockhorst:1994jf , Mirazita:1997yp , Bennhold:1997mg , SAPHIR:1998noz , SAPHIR:1998fev , Muccifora:1998ct , SAPHIR:1999wfu , Barth:2001cb , Barth:2003kv , Barth:2003bq , SAPHIR:2003lnh , Glander:2003jw , Wu:2005wf , Lawall:2005np , Wieland:2010run .
@@
The Crystal Barrel detector CrystalBarrel:1992qav had previously been used at the CERN Low Energy Antiproton Ring. The Two-Arm-Photon Spectrometer Novotny:1991ht was dismantled at Mainz and installed in Bonn to provide better granularity for the forward hole of the Crystal Barrel detector. Data were collected on meson photoproduction off protons or neutrons, both without polarization and with various combinations of photon and target polarizations CB-ELSA:2003rxy , CB-ELSA:2004sqg , CBELSATAPS:2005iwc , CB-ELSA:2007htf , CB-ELSA:2007wep , CBELSATAPS:2007oqn , CBELSA:2007vce , Thoma:2007bm , Sarantsev:2007aa , CB-ELSA:2007xbv , CBELSA:2008wwq , CBELSATAPS:2019ylw , CBELSATAPS:2020cwk , CBELSATAPS:2021osa , CBELSATAPS:2022uad .
[…]
@@
The Thomas Jefferson National Accelerator Facility Laboratory (JLab) in Newport News, Virginia, delivered its first beam in 1994. A central topic in its wide program is pursued with the CLAS detector CLAS:2003umf , measuring a variety of meson final states using unpolarized, linearly polarized, and circularly polarized photon beams with energies up to 5.6 GeV. Results are documented in CLAS:2023ddn , CLAS:2021jhm , CLAS:2021hex , CLAS:2021osv , CLAS:2020ngl , CLAS:2020rdz , CLAS:2020spy , Mokeev:2020hhu , CLAS:2018mmb , CLAS:2018kvn , CLAS:2018azo , CLAS:2018avi , CLAS:2018drk , CLAS:2018gxz , CLAS:2020rdz , CLAS:2018xbd , Anisovich:2018yoo , CLAS:2017kyf , CLAS:2017vxx , CLAS:2017jrx , CLAS:2017sgi , Anisovich:2017pox , Anisovich:2015gia , Anisovich:2017ygb , Anisovich:2017xqg , CLAS:2017gsu , CLAS:2017dco , CLAS:2017kua , CLAS:2017yjv , CLAS:2017rxe , CLAS:2016zjy , CLAS:2016wrl , CLAS:2015pjm , CLAS:2015ykk , CLAS:2014tbc , Anisovich:2014yza , CLAS:2013pcs , CLAS:2013jlg , CLAS:2013rxx , CLAS:2013qgi , CLAS:2013rxx , CLAS:2013owj , Anisovich:2013jya , CLAS:2013rjt , Afanasev:2012fh , Price:2004hr , CLAS:2011aa , CLAS:2010gcv , CLAS:2010aen , Qian:2010rr , CLAS:2009rdi , CLAS:2009ngd , CLAS:2009tyz , CLAS:2009wde , CLAS:2009hpc , Chen:2009sda , CLAS:2008ycy , Dugger:2007bt , CLAS:2007xhu , Guo:2007dw , CLAS:2007kab , CLAS:2006pde , CLAS:2005koo , CLAS:2006czw , CLAS:2006rru , CLAS:2006anj , CLAS:2005rxe , CLAS:2005oqk , CLAS:2005lui , Burkert:2005ft , CLAS:2004gjf , CLAS:2003wfm , CLAS:2003zrd , CLAS:2003yuj , CLAS:2002cdi , CLAS:2002rxi , CLAS:2001zxv , CLAS:2000kid , Manak:2000ty , JeffersonLabE94014:1998czy , Frolov:1998pw .
@@
The GRAAL Collaboration scattered a beam of ultraviolet laser light off the electron beam in the straight section of the European Synchrotron Radiation Facility, thus producing narrowly collimated $\gamma$-rays with a maximum energy of 1.47 GeV. The energy distribution is flat, and the $\gamma$-rays carry the polarization of the primary photons. GRAAL results can be found in Refs. GRAAL:2000gnk , GRAAL:2000qng , GRAAL:2002mil , Assafiri:2003mv , Gurzadyan:2004rx , GRAAL:2005mor , Lleres:2007tx , GRAAL:2007gsc , Bartalini:2008zza , GRAAL:2013tzy , LeviSandri:2014uhc . The BGO calorimeter is a central part of the GRAAL detector. After the end of data taking in Grenoble in 2008, the calorimeter was transferred to Bonn and became the central component of the BGO-OD experiment Schmieden:2018yjy , BGO-OD:2019utx at ELSA. Results of the BGO-OD experiment at Bonn can be found in Refs. Jude:2020byj , Alef:2020yul , BGOOD:2021sog , BGOOD:2021oxp , Jude:2022atd , Figueiredo:2024zmh , Rosanowski:2024rww .
@@
At the 8 GeV synchrotron radiation facility Spring-8, situated in Harima Science Park, Hyōgo, an ultraviolet photon beam is backscattered off the circulating electron beam, producing collimated $\gamma$-rays up to 2.4 GeV Nakano:2000ku , Nakano:2001xp . Charged particles produced in $\gamma N$ reactions in the forward direction are momentum-analyzed in a dipole magnet, with particle identification provided by time-of-flight and Cherenkov counters. The facility started operating in 1999. Later, a new experiment was installed with an egg-shaped BGO calorimeter, forward drift chambers, and a time-of-flight wall Muramatsu:2021bpl . Physics results obtained by the LEPS and LEPS2/BGOegg collaborations have been reported in Refs. Matsumura:2003mw , LEPS:2003wug , LEPS:2003buk , CDF:2003epb , Nakano:2004cr , Ishikawa:2004id , LEPS:2005hax , LEPS:2005hji , Kohri:2006yx , Chang:2007fc , LEPS:2007nri , Sumihama:2007qa , Sumihama:2007gfo , Miwa:2007xk , Niiyama:2008rt , LEPS:2008ghm , LEPS:2008azb , Muramatsu:2009zp , LEPS:2009isz , LEPS:2009nuw , LEPS:2009pib , LEPS:2010ovn , Hwang:2012zza , Hwang:2013usa , Morino:2013raa , LEPS:2013dqu , LEPS:2016ljn , LEPS:2017jqw , LEPS:2017vas , LEPS:2017nqz , LEPS:2017pzl , LEPS:2018pbi , LEPS2:2019bek , LEPS2BGOegg:2020cth , LEPS2:2020ttk , LEPS2BGOegg:2022dop , LEPS2BGOegg:2023ssr .